On June 30, 2017, Wisconsin, Inc., issued $267,350 in debt and 18,400 new shares of its $10 par value stock to Badger Company owners in exchange for all of the outstanding shares of that company. Wisconsin shares had a fair value of $40 per share. Prior to the combination, the financial statements for Wisconsin and Badger for the six-month period ending June 30, 2017, were as follows: Wisconsin Badger Revenues $ (985,000 ) $ (339,000 ) Expenses 720,000 201,000 Net income $ (265,000 ) $ (138,000 ) Retained earnings, 1/1 $ (843,000 ) $ (208,000 ) Net income (265,000 ) (138,000 ) Dividends declared 106,250 0 Retained earnings, 6/30 $ (1,001,750 ) $ (346,000 ) Cash $ 110,750 $ 59,000 Receivables and inventory 433,000 180,000 Patented technology (net) 929,000 372,000 Equipment (net) 727,000 619,000 Total assets $ 2,199,750 $ 1,230,000 Liabilities $ (568,000 ) $ (414,000 ) Common stock (360,000 ) (200,000 ) Additional paid-in capital (270,000 ) (270,000 ) Retained earnings (1,001,750 ) (346,000 ) Total liabilities and equities $ (2,199,750 ) $ (1,230,000 ) Wisconsin also paid $30,800 to a broker for arranging the transaction. In addition, Wisconsin paid $43,100 in stock issuance costs. Badger’s equipment was actually worth $765,250, but its patented technology was valued at only $350,700. What are the consolidated balances for the following accounts? (Input all amounts as positive values)
Answer #1

ANSWER :-

Net income =265000(Wisconsin) - 30800 234,200
Retained earnings 1.1.17 843000

Patented technology =929000+350700

1,279,700

Goodwill (see working note below)

Liabilities =568000+414000+267350(debt)

1,249,350

Common stock =360000+(18400*10)

544,000

Add: paid-in capital =27000+(18400*(40-10))-43100(issuance cost)

535,900

Working note:Calculation of Goodwill

Consideration transfered (Face value ) for Badger stock =(18400*40) +267350(debt)

1,003,350

Less:Book value of Badger=1230000(total assets -414000(liabilities)

816,000

Fair value in excess of book value

187,350

Excess fair value (under valued equipment) =765250-619000

146,250

Excess fair value (over valued Patented technology) =372000-350700

(21,300)

Goodwill =187350-146250-(-)21300

62,400

_____________________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU**************